# Computing Chebyshev polynomials using the complex Remez algorithm

### v2

@@ 1.3 Outline
In Section 2 a short discussion concerning Tang’s algorithm from [45] is presented. In particular its relation to the computation of Chebyshev polynomials is exemplified. This section serves as the method part of the article. A psuedo-code implementation is provided in the appendix as Algorithms 1 and 2.
In Section 3 we present numerical findings related to computations of Chebyshev polynomials using Tang’s algorithm. In particular Widom factors and zeros are computed for regular polygons, the $m$-cusped hypocyloid, circular lunes and the Bernoulli lemniscate. We also compare the difference between Chebyshev polynomials and Faber polynomials for such sets.
In Section 4 the results from Section 3 are discussed and we form conjectures based on these. Our main hypothesis is that the asymptotic behavior of Faber polynomials and Chebyshev polynomials have strong ties when it comes to asymptotic zero distributions, however, when it comes to norm behavior these can behave rather differently.
2 Numerical computations of Chebyshev polynomials
In the following we consider the procedure of approximating complex-valued functions on a compact subset of the complex plane, henceforth denoted $\mathsf{E}$. Conforming to the situation considered in [45, 46] we restrict ourselves to the consideration of real linear spaces in the sense that all scalars appearing in linear combinations will be real-valued. Since any $k$-dimensional complex space can be regarded as a $2k$-dimensional space over the real numbers this is no restriction. We introduce the notation $\mathcal{C}_{\mathbb{R}}(\mathsf{E})$ to denote the linear space of complex-valued continuous functions on $\mathsf{E}$ with real linear combinations. We further let $V$ denote an $n$-dimensional subspace of $\mathcal{C}_{\mathbb{R}}(\mathsf{E})$ with an associated basis $\{\varphi_{k}\}_{k=1}^{n}$. The algorithm developed by Tang computes the best approximation $\varphi^{\ast}$ to $f$ among all elements of $V$. In other words
@@ 2 Numerical computations of Chebyshev polynomials
for every $\varphi\in V$. We assume throughout that $\varphi^{\ast}$ is unique. This will be the case when studying Chebyshev polynomials on a continuum, that is, a compact connected set containing infinitely many points. To conform to the case of Chebyshev polynomials we would let $f(z)=z^{n}$ and $\varphi$ denote a complex polynomial over $\mathbb{R}$ of degree at most $n-1$.
As usual, we let $\mathcal{C}_{\mathbb{R}}(\mathsf{E})^{\ast}$ denote the dual space of $\mathcal{C}_{\mathbb{R}}(\mathsf{E})$ and $V^{\perp}$ those linear functionals in $\mathcal{C}_{\mathbb{R}}(\mathsf{E})^{\ast}$ that vanish on $V$. Riesz’ representation theorem states that any real linear functional in $\mathcal{C}_{\mathbb{R}}(\mathsf{E})^{\ast}$ can be represented through the formula
@@
where $\mu$ is a complex Borel measure. The extension theorem of Hahn–Banach implies an elementary relation between linear functionals and distance minimizing elements for Banach spaces. From [30, Theorem 7 in §8.2] we see that
$\min_{\varphi\in V}\|f-\varphi\|_{\mathsf{E}}=\max_{\begin{subarray}{c}L\in V^{\perp}\\ \|L\|\leq 1\end{subarray}}|Lf|.$ (12)
As stated, (12) provides no substantial information on the actual maximizing linear functional. The space of all complex Borel measures on $\mathsf{E}$ may prove too unwieldly to deal with in any practical situation. However, there exists maximizing linear functionals satisfying (12) with a specific simple form as was shown by Zuhovickiĭ and Remez, see e.g. [41, Theorem 2, p. 437]. The value in (12) coincides with the maximal value of all expressions of the form
$L_{\boldsymbol{r},\boldsymbol{\alpha},\boldsymbol{z}}(f)=\sum_{j=1}^{n+1}r_{j}\mathrm{Re}(e^{-i\alpha_{j}}f(z_{j}))$ (13)
where $\boldsymbol{r}=\{r_{j}\}_{j=1}^{n+1}\in[0,1]^{n+1}$, $\boldsymbol{\alpha}=\{\alpha_{j}\}_{j=1}^{n+1}\in[0,2\pi)^{n+1}$ and $\boldsymbol{z}=\{z_{j}\}_{j=1}^{n+1}\in\mathsf{E}^{n+1}$ are subject to the constraints:
$\sum_{j=1}^{n+1}r_{j}=1,$ (14)
$L_{\boldsymbol{r},\boldsymbol{\alpha},\boldsymbol{z}}(\varphi)=\sum_{j=1}^{n+1}r_{j}\mathrm{Re}(e^{-i\alpha_{j}}\varphi(z_{j}))=0,\quad\forall\varphi\in V.$ (15)
The goal of using Tang’s algorithm, which is further illustrated in Appendix A, concerns the computation of the maximizing functional. The algorithm produces a sequence of linear functionals $\{L_{\boldsymbol{r}^{(\nu)},\boldsymbol{\alpha}^{(\nu)},\boldsymbol{z}^{(\nu)}}\}$ together with an associated sequence of approximants $\{\varphi^{(\nu)}\}$ that satisfy that $L_{\boldsymbol{r}^{(\nu)},\boldsymbol{\alpha}^{(\nu)},\boldsymbol{z}^{(\nu)}}(f)$ is increasing in $\nu$ and
$L_{\boldsymbol{r}^{(\nu)},\boldsymbol{\alpha}^{(\nu)},\boldsymbol{z}^{(\nu)}}(f)\leq\|f-\varphi^{\ast}\|_{\mathsf{E}}\leq\|f-\varphi^{(\nu)}\|_{\mathsf{E}}.$ (16)
One of the novelties with Tang’s algorithm in comparison to previous algorithms at the time of its inception is that it can be shown to converge quadratically if certain conditions are met, see [45, 46] for further details. If one assumes that $\boldsymbol{r}^{(\nu)}>0$ for all sufficiently large $\nu\in\mathbb{N}$ then
@@
A simple proof of this can be found in [16]. As a consequence it follows that at least a subsequence of $\varphi^{(\nu)}$ converges to $\varphi^{\ast}$ under the assumption that the minimizer is unique. It should be mentioned that in our computation of Chebyshev polynomials, we have typically observed rapid convergence.
3 Computations of Chebyshev polynomials
We now turn to the computation of Chebyshev polynomials in the complex plane. We stress the fact that this section will only contain computational results and the discussion of these are postponed to Section 4. To translate the notation from Section 2 to the present situation we let $n$ be a specified degree and $\gamma:[0,1]\rightarrow\mathbb{C}$ a parametrization of a curve denoted $\mathsf{E}$. In order to compute $T_{n}^{\mathsf{E}}$ we let $f(t)=\gamma(t)^{n}$ and in the general case, we choose the basis as
@@ 3 Computations of Chebyshev polynomials
$=\begin{bmatrix}1&\gamma(t)&\cdots&\gamma(t)^{n-1}&i&i\gamma(t)&\cdots i\gamma(t)^{n-1}\end{bmatrix}.$
The algorithm, applied to this setting, will produce coefficients $\lambda_{1},\cdots\lambda_{2n}$ so that
@@
In many cases it is possible to exploit the symmetry of a set to reduce the size of the basis which significantly helps with speeding up the computation. As an example if $\mathsf{E}$ is conjugate symmetric meaning that
$z\in\mathsf{E}\quad\Leftrightarrow\quad\overline{z}\in\mathsf{E}$
then by the uniqueness of $T_{n}^{\mathsf{E}}$ all coefficients appearing must be real. Hence the basis can be chosen to be the $n$-dimensional real linear space spanned by
@@
In general, we have the following lemma, see also [12, Example 4.1].
Lemma 1.
Let $\mathsf{E}$ denote a compact infinite set, satisfying
$\mathsf{E}=\{e^{2\pi im/n}z:z\in\mathsf{E}\}.$
For $n\in\mathbb{N}$ and $l\in\{0,1\dotsc,m-1\}$,
$T_{nm+l}^{\mathsf{E}}(z)=z^{nm+l}+\sum_{k=0}^{n-1}a_{k}z^{km+l}=z^{l}Q_{n}(z^{m})$
@@
$e^{-2\pi iml/n}T_{nm+l}^{\mathsf{E}}(e^{2\pi im/n}z)=z^{nm+l}+\text{lower order terms},$
we see that this is a monic polynomial with the same norm as $T_{nm+l}^{\mathsf{E}}$ on $\mathsf{E}$. From uniqueness of the corresponding Chebyshev polynomial we conclude that
$e^{-2\pi iml/n}T_{nm+l}^{\mathsf{E}}(e^{2\pi im/n}z)=T_{nm+l}^{\mathsf{E}}(z)$
which immediately implies the result. ∎
As a consequence of Lemma 1 it is possible to exploit the symmetry of the underlying set in order to make further reductions on the size of the basis used in Tang’s algorithm.
We will consider the computation of Chebyshev polynomials corresponding to a plethora of sets for which the asymptotics remain unknown. Firstly we will consider the computation of Widom factors, $\mathcal{W}_{n}$ as defined in (9). Secondly we will investigate a possible connection between Chebyshev polynomials and Faber polynomials using numerical experiments. Finally we will consider the computation of zeros of $T_{n}^{\mathsf{E}}$.
Remark 1.
Let us heavily emphasize the fact that the computations performed here will provide $n$th degree monic polynomials $P_{n}$ such that $\|P_{n}\|_{\mathsf{E}}$ is close to the theoretical minimum $\|T_{n}^{\mathsf{E}}\|_{\mathsf{E}}$. Furthermore, $\|P_{n}\|_{\mathsf{E}}-\|T_{n}^{\mathsf{E}}\|_{\mathsf{E}}$ can be explicitly upper bounded in the computations using (16). This implies that Widom factors can be accurately estimated. Regarding intricate polynomial properties such as their coefficients and zeros, the algorithm has to be used with care. Although it is true that if $\|P\|_{\mathsf{E}}$ is close to $\|T_{n}^{\mathsf{E}}\|_{\mathsf{E}}$ then their distance is small in every measurable way, it is in general difficult to quantify this. We remark however, that the computations are consistent in the sense that the behaviors here exhibited do not change as the precision is increased further.
3.1 Computations of Widom Factors
As was already stated in Section 1 we recall that if $\mathsf{E}$ denotes the closure of a Jordan domain with $C^{1+\alpha}$ boundary then it is known that $\mathcal{W}_{n}(\mathsf{E})\rightarrow 1$ as $n\rightarrow\infty$, see [42, 43, 51]. If $\mathsf{E}$ is convex it is possible to conclude that $\mathcal{W}_{n}(\mathsf{E})\leq 2$, see [28]. If $\mathsf{E}$ is a quasi-disk then $\mathcal{W}_{n}(\mathsf{E})$ is known to be bounded [4]. Likewise, the assumption that the outer boundary of $\mathsf{E}$ consists of dini-smooth arcs which are disjoint apart from their endpoints which do not have external cusps also implies that $\mathcal{W}_{n}(\mathsf{E})$ is bounded, see [49, Theorem 2.1]. Informally stated, an external cusp is a point where the intersecting arcs form an angle of $2\pi$ on the interior of $\mathsf{E}$ so that it “points away” from the unbounded complement. Apart from these results, very few general estimates exist regarding Widom factors related to compact sets, even with the additional assumption that they are closures of Jordan domains.
We stress the fact that $\mathcal{W}_{n}(\mathsf{E})$ is invariant under dilations and translations in the sense that for any $\alpha,\beta\in\mathbb{C}$ with $\alpha\neq 0$ we have
@@ 3.1 Computations of Widom Factors
see [18]. Therefore it is always possible to rotate and scale the set in question in a way so that symmetries can be easily exploited without affecting the Widom factors. We remind the reader that in the following section we will simply present the results of numerical computations and leave the discussion of these results to Section 4.
3.1.1 Regular polygon
Simple examples of piecewise analytic Jordan domains with corners are the regular polygons or simply $m$-gons if they have $m$ sides of equal length. Due to the convexity of such sets we immediately gather that if $\mathsf{E}$ is a regular polygon then $\mathcal{W}_{n}(\mathsf{E})\leq 2$. It is not known whether the sequence $\{\mathcal{W}_{n}(\mathsf{E})\}$ converges in this case and we therefore proceed with studying the corresponding Widom factors numerically. Previous numerical considerations for Chebyshev polynomials corresponding to a square have been undertaken in [32] for degrees up to 16. These, however, lack the perspective of Widom factors. The logarithmic capacity of a regular $m$-gon $\mathsf{E}$ can be found in [33, Table 5.1]. It is there stated that
$\operatorname{Cap}(\mathsf{E})=\frac{\Gamma(1/m)}{2^{1+2/m}\pi^{1/2}\Gamma(1/2+1/m)}\cdot\text{side length}(\mathsf{E}).$ (17)

### v3

@@ 1.3. Outline
In Section 2, a short discussion concerning Tang’s algorithm from [50] is presented. In particular, we demonstrate how it can be used to compute Chebyshev polynomials. This section serves as the method part of the article. A pseudocode implementation is provided in the appendix as Algorithm 1.
In Section 3, we present numerical findings related to computations of Chebyshev polynomials using Tang’s algorithm. In particular, Widom factors and zeros are computed for regular polygons, the $m$-cusped hypocycloid, circular lunes and the Bernoulli lemniscate. We also compare the difference between Chebyshev polynomials and Faber polynomials for some of these sets.
In Section 4, the results from Section 3 are discussed and we formulate hypotheses about the behavior of Chebyshev polynomials based on these numerical experiments. Our main hypothesis is that the asymptotic behavior of Faber polynomials and Chebyshev polynomials is closely related.
Acknowledgement
The authors are grateful to Frank Wikström and Jacob Stordal Christiansen, both at Lund University. The former for providing excellent advice on the presentation of the numerical experiments in this manuscript and the latter for helpful discussions on the material presented herein. The anonymous referee is also thanked for constructive comments.
The research of L. Hübner was supported by Odysseus grant number G0DDD23N from the FWO.
2. Numerical computations of Chebyshev polynomials
In the following, we consider the procedure of approximating complex-valued functions on a connected compact subset of the complex plane, henceforth denoted $\mathsf{E}$. In line with the setting considered in [50, 51], we restrict ourselves to real linear spaces in the sense that all scalars appearing in linear combinations will be real-valued. Since any $k$-dimensional complex space can be regarded as a $2k$-dimensional space over the real numbers this does not impose any restriction. We introduce the notation $\mathcal{C}_{\mathbb{R}}(\mathsf{E})$ to denote the linear space of complex-valued continuous functions on $\mathsf{E}$, where linear combinations are formed with real scalars. We further let $V$ denote an $n$-dimensional subspace of $\mathcal{C}_{\mathbb{R}}(\mathsf{E})$ with an associated basis $\{\varphi_{k}\}_{k=1}^{n}$. The algorithm developed by Tang computes the best approximation $\varphi^{\ast}$ to $f$ in the space $V$. In other words,
@@ 2. Numerical computations of Chebyshev polynomials
for every $\varphi\in V$. We assume throughout that $\varphi^{\ast}$ is unique. This will be the case when studying Chebyshev polynomials on a continuum, that is, a compact connected set containing infinitely many points. To conform to the case of Chebyshev polynomials, we let $f(z)=z^{n}$ and $\varphi$ denote a complex polynomial of degree at most $n-1$.
As usual, we let $\mathcal{C}_{\mathbb{R}}(\mathsf{E})^{\ast}$ denote the dual space of $\mathcal{C}_{\mathbb{R}}(\mathsf{E})$ and $V^{\perp}$ those linear functionals in $\mathcal{C}_{\mathbb{R}}(\mathsf{E})^{\ast}$ that vanish on $V$. The Riesz representation theorem states that any real linear functional in $\mathcal{C}_{\mathbb{R}}(\mathsf{E})^{\ast}$ can be represented through the formula
@@
where $\mu$ is a complex Borel measure. The extension theorem of Hahn–Banach implies an elementary relation between linear functionals and distance minimizing elements for Banach spaces. From [34, Theorem 7 in §8.2] we see that
(12) $\min_{\varphi\in V}\|f-\varphi\|_{\mathsf{E}}=\max_{\begin{subarray}{c}L\in V^{\perp}\\ \|L\|\leq 1\end{subarray}}|Lf|.$
As stated, (12) provides no substantial information on the actual maximizing linear functional. The space of all complex Borel measures on $\mathsf{E}$ may prove too unwieldy to deal with in any practical situation. However, there exists maximizing linear functionals satisfying (12) with a specific simple form as was shown by Zuhovickiĭ and Remez, see e.g. [46, Theorem 2, p. 437]. The value in (12) coincides with the maximal value of all expressions of the form
(13) $L_{\boldsymbol{r},\boldsymbol{\alpha},\boldsymbol{z}}(f)=\sum_{j=1}^{n+1}r_{j}\mathrm{Re}(e^{-i\alpha_{j}}f(z_{j}))$
where $\boldsymbol{r}=\{r_{j}\}_{j=1}^{n+1}\in[0,1]^{n+1}$, $\boldsymbol{\alpha}=\{\alpha_{j}\}_{j=1}^{n+1}\in[0,2\pi)^{n+1}$ and $\boldsymbol{z}=\{z_{j}\}_{j=1}^{n+1}\in\mathsf{E}^{n+1}$ are subject to the following constraints:
(14) $\sum_{j=1}^{n+1}r_{j}=1,$
(15) $L_{\boldsymbol{r},\boldsymbol{\alpha},\boldsymbol{z}}(\varphi)=\sum_{j=1}^{n+1}r_{j}\mathrm{Re}(e^{-i\alpha_{j}}\varphi(z_{j}))=0,\quad\forall\varphi\in V.$
The goal of using Tang’s algorithm, which is further illustrated in Appendix A, is to compute the maximizing functional as in (13). The algorithm produces a sequence of linear functionals $\{L_{\boldsymbol{r}^{(\nu)},\boldsymbol{\alpha}^{(\nu)},\boldsymbol{z}^{(\nu)}}\}$ together with an associated sequence of approximants $\{\varphi^{(\nu)}\}$, such that $L_{\boldsymbol{r}^{(\nu)},\boldsymbol{\alpha}^{(\nu)},\boldsymbol{z}^{(\nu)}}(f)$ is increasing in $\nu$ and
(16) $L_{\boldsymbol{r}^{(\nu)},\boldsymbol{\alpha}^{(\nu)},\boldsymbol{z}^{(\nu)}}(f)\leq\|f-\varphi^{\ast}\|_{\mathsf{E}}\leq\|f-\varphi^{(\nu)}\|_{\mathsf{E}}.$
Tang’s algorithm will be applied until a wanted accuracy level, in terms of the relative error given by
(17) $\frac{\|f-\varphi^{(\nu)}\|_{\mathsf{E}}-L_{\boldsymbol{r}^{(\nu)},\boldsymbol{\alpha}^{(\nu)},\boldsymbol{z}^{(\nu)}}(f)}{L_{\boldsymbol{r}^{(\nu)},\boldsymbol{\alpha}^{(\nu)},\boldsymbol{z}^{(\nu)}}(f)}$
is reached. Throughout, any reference to the precision of our computations will be in terms of this relative error.
One of the novel features of Tang’s algorithm in comparison to previous algorithms at the time of its inception is that it can be shown to converge quadratically if certain conditions are met, see [50, 51] for further details. Assuming that $\boldsymbol{r}^{(\nu)}>0$ for all sufficiently large $\nu\in\mathbb{N}$ then
@@
A simple proof of this can be found in [17]. We note that in our computations of Chebyshev polynomials, rapid convergence is typically observed.
3. Computations of Chebyshev polynomials
We now turn to the computation of Chebyshev polynomials in the complex plane and stress the fact that this section will only contain computational results. These results will be discussed in Section 4. Let $\mathsf{E}$ be a polynomially convex set in the complex plane, by the Maximum principle $T_{n}^{\mathsf{E}}=T_{n}^{\partial\mathsf{E}}$. To translate the notation from Section 2, we let $n$ be a specified degree and $\gamma:[0,1]\rightarrow\mathbb{C}$ a parametrization of $\partial\mathsf{E}$. In order to compute $T_{n}^{\mathsf{E}}$ we let $f(t)=\gamma(t)^{n}$ and in the general case, we choose the basis as
@@ 3. Computations of Chebyshev polynomials
$=\begin{bmatrix}1&\gamma(t)&\cdots&\gamma(t)^{n-1}&i&i\gamma(t)&\cdots&i\gamma(t)^{n-1}\end{bmatrix}.$
The algorithm, applied to this setting, is designed to produce coefficients $\lambda_{1},\dots,\,\lambda_{2n}$ such that
@@
In many cases it is possible to exploit the symmetry of a set to reduce the size of the basis which significantly helps with speeding up the computation. As an example, if $\mathsf{E}$ is symmetric with respect to complex conjugation, meaning that
$z\in\mathsf{E}\quad\Leftrightarrow\quad\overline{z}\in\mathsf{E},$
then by the uniqueness of $T_{n}^{\mathsf{E}}$ all coefficients must be real. Hence the basis can be chosen to be the $n$-dimensional real linear space spanned by
@@
Using the following lemma, it is possible to exploit the symmetry of the underlying set in order to make further reductions on the size of the basis used in Tang’s algorithm, see also [13, Example 4.1].
Lemma 3.1.
Let $\mathsf{E}$ denote a compact infinite set satisfying
(18) $\mathsf{E}=\{e^{2\pi i/m}z:z\in\mathsf{E}\}$
for some $m\in\mathbb{N}$. If $n\in\mathbb{N}$ and $l\in\{0,1\dotsc,m-1\}$ then
(19) $T_{nm+l}^{\mathsf{E}}(z)=z^{nm+l}+\sum_{k=0}^{n-1}a_{k}z^{km+l}=z^{l}Q_{n}(z^{m})$
@@
$e^{-2\pi il/m}T_{nm+l}^{\mathsf{E}}(e^{2\pi i/m}z)=z^{nm+l}+\text{lower order terms},$
we see that this is a monic polynomial and from (18) we conclude that it has the same norm as $T_{nm+l}^{\mathsf{E}}$ on $\mathsf{E}$. From the uniqueness of the corresponding Chebyshev polynomial
$e^{-2\pi il/m}T_{nm+l}^{\mathsf{E}}(e^{2\pi i/m}z)=T_{nm+l}^{\mathsf{E}}(z).$
The coefficient in front of $z^{j}$ on the left-hand side is equal to $e^{2\pi i(j-l)/m}$ multiplied with the coefficient in front of $z^{j}$ on the right-hand side. Hence, if $j-l$ fails to be a multiple of $m$ then this coefficient must be $0$ This establishes that (19) holds. ∎
We will consider the computation of Chebyshev polynomials corresponding to a variety of sets for which their asymptotic behavior remains unknown. First, we compute the Widom factors $\mathcal{W}_{n}$. Next, we explore a possible connection between Chebyshev polynomials and Faber polynomials using numerical experiments. Finally, we examine the distribution of zeros of $T_{n}^{\mathsf{E}}$.
All computations are carried out in Python using the arbitrary-precision arithmetic module mpmath. For each computation, we use a working accuracy of at least 50 significant digits to avoid phenomena caused by numerical instabilities.
3.1. Computations of Widom factors
As was already stated in Section 1, recall that if $\mathsf{E}$ denotes the closure of a Jordan domain with $C^{1+\alpha}$ boundary then it is known that $\mathcal{W}_{n}(\mathsf{E})\rightarrow 1$ as $n\rightarrow\infty$, see [47, 48, 57]. If $\mathsf{E}$ is convex it is possible to conclude that $\mathcal{W}_{n}(\mathsf{E})\leq 2$, see [42]. If $\mathsf{E}$ is a quasi-disk then $\mathcal{W}_{n}(\mathsf{E})$ is known to be bounded [5]. Likewise, if the outer boundary of $\mathsf{E}$ consists of Dini-smooth arcs which are disjoint apart from their endpoints and do not form inward pointing cusps then the sequence $\{\mathcal{W}_{n}(\mathsf{E})\}$ is bounded, see [55, Theorem 2.1]. Informally stated, an inward pointing cusp (also external cusp in the literature) is a point where the intersecting arcs form an angle of $2\pi$ within the interior of $\mathsf{E}$ so that it “points away” from the unbounded complement of $\mathsf{E}$. Apart from these results, very few general estimates exist regarding Widom factors for compact sets, even with the additional assumption that they are closures of Jordan domains.
Note that $\mathcal{W}_{n}(\mathsf{E})$ is invariant under dilations and translations in the sense that for any $\alpha,\beta\in\mathbb{C}$ with $\alpha\neq 0$ we have
@@ 3.1. Computations of Widom factors
see [20]. Thus, the set may always be rotated and scaled so that symmetries can be easily exploited without affecting the Widom factors. The results of the numerical computations are presented below, while their discussion is deferred to Section 4.
3.1.1. Regular polygons
Simple examples of piecewise-analytic Jordan domains with corners are the regular polygons (also called $m$-gons) if they have $m$ sides of equal length. Due to the convexity of such sets we immediately gather that if $\mathsf{E}$ is a regular polygon then $\mathcal{W}_{n}(\mathsf{E})\leq 2$. It is not known whether the sequence $\{\mathcal{W}_{n}(\mathsf{E})\}$ converges in this case and we thus study the corresponding Widom factors numerically. Previous numerical studies for Chebyshev polynomials corresponding to a square have been conducted in [37] for degrees up to 16. These, however, lack the perspective of Widom factors. The logarithmic capacity of a regular $m$-gon $\mathsf{E}$ can be found in [38, Table 5.1]. It is stated there that
(20) $\operatorname{Cap}(\mathsf{E})=\frac{\Gamma(1/m)}{2^{1+2/m}\pi^{1/2}\Gamma(1/2+1/m)}\cdot\text{side length}(\mathsf{E}).$